OpenStudy (anonymous):

Trig - solve for x if 0deg < x < 360deg 4 sin x - 2 csc x = 0

OpenStudy (amistre64):

4sin(x) -2csc(x) = 0 4sin(x) = 2csc(x) 2sin(x) = csc(x)

OpenStudy (amistre64):

2sin(x) = 1/sin(x) 2sin^2(x) = 1 sin^2(x) = 1/2

OpenStudy (amistre64):

sin(x) = 1/sqrt(2) x = 45 degrees or pi/4 in quadrant 1 and x = 135 degrees or 3pi/4 in quadrant 2

OpenStudy (anonymous):

Wow, thanks a whole bunch!!!

OpenStudy (amistre64):

your welcome...